OpenStudy (anonymous):

Find the difference. 4ab2 - 3ab2 Select one of the options below as your answer: A. 1 B. a C. ab2 D. ab4

OpenStudy (anonymous):

Is it (4ab)^2 - (3ab)^2 you are talking of?

OpenStudy (anonymous):

Yes ,

OpenStudy (anonymous):

(4ab)^2 - (3ab)^2 = (4ab - 3ab)(4ab + 3ab)

OpenStudy (anonymous):

I Dont Know Any Of It

OpenStudy (anonymous):

ok, I will explain now

OpenStudy (anonymous):

Okay ,

OpenStudy (anonymous):

\[a^2-b^2=(a + b)(a - b)\]

OpenStudy (anonymous):

This is what we call difference of two squares

OpenStudy (jiteshmeghwal9):

solving these like terms is too easy